Consider a titration between 40.0 mL of .200 M HNO3 with .100 M NaOH. Calculate the pH at specific points given throughout the titration.
A. No NaOH added
B. 10.0 mL NaOH
c. 40.0 mL NaOH
D. 80.0 mL NaOH
E. 100.0 mL NaOH
1
Expert's answer
2019-12-12T05:46:32-0500
"NaOH(aq)+ HNO_3(aq) =NaNO_3(aq) + H_2O(l)"

A. No NaOH is added

"c(H^+)=c(HNO3) = 0.200 M"

"pH = -log_{10}[H^+]=-log_{10}(0.200) = 0.699"


B. 10 mL of NaOH are added

"n(HNO3) = c\times V = 0.200\times0.040= 0.008 mol"

"n(NaOH) = c\times V = 0.100\times 0.010 = 0.001 mol"

According to equation mole ratio "n(NaOH):n(HNO3) = 1:1" , then moles of HNO3 used for reaction are "n(HNO3) = n(NaOH) = 0.001 mol" , moles left after titration are "n(HNO_3)_{left} = 0.008-0.001 = 0.007 mol"


"c(H^+)= \frac{n}{V} = \frac{0.007}{0.04+0.01} = 0.14 M"


"pH = -log_{10}(0.14) = 0.854"


C. 40 .0 mL NaOH are added

"n(HNO3) = 0.008 mol"

"n(NaOH) = c\times V = 0.1\times 0.04 = 0.004 mol"

According to equation mole ratio n(NaOH):n(NaOH) , then moles of HNO3 used for reaction are n(HNO3) = n(NaOH) = 0.004 mol, moles left after titration are

"n(HNO_3)_{left} = 0.008-0.004 = 0.004 mol"

"c(H^+) = \frac{n}{V} = \frac{0.004}{0.04+0.04} = 0.05 M"


"pH = -log_{10}(0.05) = 1.3"


D. 80 mL of NaOH are added

"n(HNO3) = 0.008 mol"

"n(NaOH) = c\times V = 0.1\times 0.08=0.008 mol"

According to equation mole ratio n(NaOH):n(NaOH)=1:1 , then moles of HNO3 used for reaction are n(HNO3) = n(NaOH) = 0.008 mol, moles left after titration are

"n(HNO_3)_{left} = 0.008-0.008 = 0"

pH is determined by the "H^+" ions produced by the water dissosiation

"[H^+] = 1\times 10^{-7} M"


"pH = -log_{10}(1\times10^{-7}) = 7"


E. 100 mL of NaOH are added

"n(HNO3) = 0.008 mol"

"n(NaOH) = 0.1\times 0.1 = 0.01 mol"

Moles of NaOH left:

"n(NaOH)_{left} = 0.01-0.008 = 0.002 mol"

"n(OH^-) = n(NaOH) = 0.002 mol"

"c(OH^-) = \frac{n}{V} = \frac{0.002} {0.04+0.1} = 0.0143 M"

"pOH = -log_{10}[OH^-] = -log_{10}(0.0143) = 1.845"

As pH + pOH =14, then pH = 14-1.845 = 12.15